You intend to conduct a goodness-of-fit test for a multinomial distribution with 7 categories. You collect data from 74 subje11.2
Answer #1

Question 1

You intend to conduct a goodness-of-fit test for a multinomial distribution with 7 categories. You collect data from 74 subjects.

What are the degrees of freedom for the χ2 distribution for this test?

Solution:

We are given

Number of categories = N = 7

Degrees of freedom = df = N - 1 = 7 - 1 = 6

d.f. = 6

Question 2

Solution:

Here, we have to use chi square test for goodness of fit.

Null hypothesis: H0: Data follows the given distribution.

Alternative hypothesis: Ha: Data do not follow the given distribution.

We are given

Number of categories = N = 4

Degrees of freedom = df = N - 1 = 3

Test statistic formula is given as below:

Chi square = ∑[(O – E)^2/E]

Where, O is observed frequencies and E is expected frequencies.

Calculation tables for test statistic are given as below:

Category

Prob.

O

E

(O - E)^2/E

A

0.2

35

145*0.2 = 29

1.24137931

B

0.3

30

43.5

4.189655172

C

0.3

34

43.5

2.074712644

D

0.2

46

29

9.965517241

Total

1

145

145

17.47126437

Category

Observed Frequency

Expected Frequency

A

35

29

B

30

43.5

C

34

43.5

D

46

29

Total

145

145

Test Statistic = Chi square = ∑[(O – E)^2/E] = 17.47126437

χ2 = 17.471

P-value = 0.001

(By using Chi square table or excel)